A project has the following cash flows:
Year   Cash Flows
0   -5000.0
1   2010.0
2   1820.0
3   1750.0
4   1660.0
The required return is 9.4 percent. What is the payback period for this project?
      
2.16
      
2.22
      
2.67
      
3.18
      
2.26
Answer #1

C10 fac =2+1170/1750 A B C D 1 2 year Cash flows Cumulative cash flows 3 0 $ (5,000.00) $ (5,000.00) 4 1 $ 2,010.00 $ (2,990.

C10 fo =2+1170/1750 A B С 1 2 year Cash flows |-5000 30 4 1 2010 Cumulative cash flows =B3 =C3+B4 =C4+B5 =C5+B6 52 6 3 1820 1